# Methods Paper Version 19

### v1

@@ II.2 Formal solutions
and $c$ given by Eq. (32) are $\mathbb{\color[rgb]{0,0,0}N}$-dimensional vectors.
Equations (33) can be resolved step-by-step. First, we solve this matrix equation for the first entrance channel with $i=1$ and obtain $v$ coefficients $b_{q}^{1}(E)={c}^{1}_{q}$, $q=0,...\,,v-1$, and $w$ $S$-matrix elements ${\mathbb{S}_{j1}(E)=-c^{1}_{j+v-1}}$, $j=1,...\,,w$ at any desired energy $E$. Next, we solve Eq. (33) for $i=2$. The $S$-matrix element ${\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)=-c^{1}_{v+1}}$ in the right-hand-side of Eq. (34d) was already obtained at the previous step. With the help of Eq. (33) we now obtain $v$ coefficients ${b_{q}^{2}(E)={c}^{2}_{q}}$, $q=0,...\,,v-1$, and $w-1$ $S$-matrix elements ${\mathbb{S}_{j2}(E)=-c^{2}_{j+v-2}}$, $j=2,...\,,w$. Now we have $S$-matrix elements $\mathbb{S}_{3j}(E)=-c^{j}_{v+3-j}$ with $j=1$, 2 in the right-hand-side of Eq. (33) for $i=3$ and solve it. Next we solve Eq. (33) for $i=4$, 5, etc.
@@
Following the ideas of the HORSE approach BangHORSE ; Zaitsev1998 , the potential energy operator $V$ can be approximated by a finite matrix $V^{tr}$ in the oscillator basis,
$V\approx V^{tr}=\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle\langle\phi_{ni}|V|\phi_{n^{\prime}j}\rangle\langle\phi_{n^{\prime}j}|\\ \equiv\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|,$ (40)
since the potential energy matrix elements in the oscillator basis $V_{ni,n^{\prime}j}$ decrease with $n$ and/or $n^{\prime}$ and can be neglected at large $n$ and/or $n^{\prime}$ as compared to the non-zero kinetic energy matrix elements $T_{ni,n^{\prime}j}=\langle\phi_{ni}|T|\phi_{n^{\prime}j}\rangle=T^{l_{i}}_{nn^{\prime}}\delta_{ij}$,
@@
$-\sum_{j^{\prime}=1}^{i-1}\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j^{\prime}}-1}\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}C^{+}_{n^{\prime}l_{j^{\prime}}}\!(k)\,V_{nj,n^{\prime}j^{\prime}}c^{j^{\prime}}_{v+i-j^{\prime}}-\sum_{j^{\prime}=1}^{i-1}\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{kS_{0l_{j^{\prime}}}(k)}c^{j^{\prime}}_{v+i-j^{\prime}},\quad$$\begin{cases}i>1,\\ q=0,...\,,N-i,\end{cases}$ (43d)
@@
where ¯$\bar{\mathcal{M}}_{i}=\min\{\bar{M}_{i},\mathcal{N}_{i}\}$, $\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}$ is a complex conjugate of $\bar{a}_{qn;i}^{j}$, and, as previously mentioned, we set $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for $q=0,...,{v-1}$ and arbitrarily chosen coefficients for $q=v,...,{N-i}$.
@@
Efros mentioned VDEfros that, contrary to the conventional Hulthén–Kohn variational method Kohn1948 ; Moses1953 ; Moses , his approach does not make use of “free-free” matrix elements like $\langle\eta^{\pm}_{i}|H-E|\eta^{\pm}_{j}\rangle$ but, instead, includes only “bound-bound” $\langle\bar{\beta}^{i}_{q}|{H-E}|\beta_{q^{\prime}}\rangle$ and “bound-free” $\langle\bar{\beta}^{i}_{q}|{H-E}|\eta^{\pm}_{j}\rangle$ matrix elements which are easier to calculate. We note that Eqs. (43) include matrix elements of the “bound-bound” type only, which are evaluated within a restricted Hilbert space spanned by a limited set of oscillator functions.
@@
We should also note that in many-body applications, Eqs. (43b)–(43d) become more complicated. The problem is that the terms ±$\pm\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{k\,S_{0l_{j^{\prime}}}\!(k)}$ in these equations account for the kinetic energy of relative motion of colliding particles $T^{rel}$ only. In the many-body case, the colliding particles have an internal structure. Therefore, the Hamiltonian includes the kinetic energy of the internal motion of these particles $T^{int}$, which enters Eqs. (43b)–(43d) in the same manner as potential energy terms. The separation of the total kinetic energy operator $T=T^{rel}+T^{int}$ into the relative $T^{rel}$ and internal $T^{int}$ parts may be non-trivial within NCSM but will be much more natural and simpler in approaches like the RGM.
@@
In the case of the standing-wave approach starting from the wave-function ansatz (19), one can use all the above equations of this subsection with replacements $-\mathbb{S}_{ij}(E)\rightarrow K_{ij}(E)$, $\tilde{h}^{-}_{l_{i}}(kr)\rightarrow j_{l_{i}}(kr)$, ${\tilde{h}^{+}_{l_{i}}(kr)\rightarrow\tilde{n}_{l_{i}}(kr)}$, $C^{-}_{nl_{i}}(k)\rightarrow S_{nl_{i}}(k)$, ${C^{+}_{nl_{i}}(k)\rightarrow C_{nl_{i}}(k)}$, having in mind that with these replacements, Eqs. (37) should be substituted by the following relations:
$(T-E)j_{l_{i}}(kr)\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})$ $=0,$ (37a)
$(T-E)\tilde{n}_{l_{i}}(kr)\mathscr{Y}_{JM_{J}}^{l_{i}S_{i}}(\Omega_{\hat{r}})$ $=\frac{\hbar^{2}}{2m}\frac{\phi_{ni}(\vec{r})}{k\,S_{0l_{i}}(k)}.$ (37b)
@@
As a result, Eqs. (43a)–(43b) preserve their structure with the replacement $C^{+}_{nl}(k)\rightarrow C_{nl}(k)$ while in Eqs. (43c)–(43d) the substitution ${C^{-}_{nl}(k)\rightarrow S_{nl}(k)}$ should be accompanied by removing the terms $-\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j}\big{)}^{\!*}}{kS_{0l_{j}}\!(k)}$.
@@ II.3 $S$-matrix poles
The vector $B^{i}$ at $i>1$ includes $S$-matrix elements [the components $c^{j}_{v+i-j}$ of the vector $c$, see Eqs. (34d) and (31)] which have poles; thus searching for poles of the vector $c^{i}$ by locating the zeroes of $\det A^{i}$ with $i>1$ is not justified and one can omit some of the $S$-matrix poles. Locating the zeroes of $\det A^{1}(E)$ in the complex energy plane can be efficiently performed by the technique of calculating contour integrals described in Ref. AMShirokovHORSE2018 .
@@
Note, Eq. (46) guarantees that Eq. (48) has a non-zero solution. However, the vector of unknowns $c^{1}$ is obtained up to an arbitrary multiplier, and the bound state wave function $u_{i}(E;\vec{r})$ should be additionally normalized. Moreover, the bound state wave function in this case is expressed as an infinite converging series of oscillator functions since the function $\eta^{+}_{i}(\vec{r})$ entering $u_{i}(E;\vec{r})$ is defined through infinite oscillator function series, see Eqs. (7) and (15).
@@ III Single-Channel Scattering
We proceed to discuss the efficiency of the Efros approach and its advantages and disadvantages using two-body single-channel scattering. As the first example, we consider $np$ scattering in the ${{}^{1}S_{0}}$ channel using the Daejeon16 $NN$ interaction Daejeon16 . In the single-channel case, it is convenient to utilize the standing-wave approach with real-valued wave functions where the only $K$-matrix element in Eq. (19) is expressed through the scattering phase shifts $\delta_{l}(E)$:
@@ III.1 Naive application
$H^{tr}=\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle\langle\phi_{n}|H|\phi_{n^{\prime}}\rangle\langle\phi_{n^{\prime}}|\equiv\sum_{n,n^{\prime}=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}|\phi_{n}\rangle H_{nn^{\prime}}\langle\phi_{n^{\prime}}|,$ (50)
where $\mathcal{\color[rgb]{0,0,0}N}$ is the number of harmonic oscillator basis functions used in the truncated wave function expansion. The truncation corresponds conventionally to the maximal value of (excitation) oscillator quanta $N_{\max}$. Having in mind future many-body applications combining the NCSM and the Efros method, we start with a Hamiltonian of this type naively supposing that solutions are obtainable within the Efros approach.
In the case of single-channel two-body scattering, the relative motion oscillator functions $\phi_{n}(\vec{r})$ are given by Eq. (3) and $2n+l\leq N_{\textrm{max}}$, where $n=0,1,...\,,\mathcal{\color[rgb]{0,0,0}N}-1$. Equation (12) takes the form
@@
$d_{n}(E)$ =$=\begin{cases}g_{n}(E)+a^{as}_{n}(E),&n=0,...\,,\mathcal{\color[rgb]{0,0,0}N}-1,\\ a^{as}_{n}(E),&n=\mathcal{\color[rgb]{0,0,0}N},...\,,\infty,\end{cases}$ (52)
$a^{as}_{n}(E)$ $=S_{nl}(k)+C_{nl}(k)\tan\delta_{l}(E),\quad n=0,...\,,\infty,$ (53)
and $\mathcal{\color[rgb]{0,0,0}N}$ energy-dependent coefficients $g_{n}(E)$ should be found within the approach. In our case of $NN$ scattering in the ${}^{1}S_{0}$ partial wave, the fixed quantum numbers are $l=0$, $S=0$, $J=0$, and $M_{J}=0$; therefore ${\mathcal{\color[rgb]{0,0,0}N}=\frac{1}{2}N_{\max}+1}$.
We need to specify the sets of SRFs $\{\beta_{q}\}$ and ${\{\bar{\beta}_{q}\}\equiv\{\bar{\beta}^{1}_{q}\}}$ entering Eqs. (34). A natural option is to use $N\leq\mathcal{\color[rgb]{0,0,0}N}$ eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the truncated Hamiltonian (50),
$\bar{\beta}_{q}(\vec{r})$ $=\beta_{q}^{\textrm{Efn}}(\vec{r}),\qquad q=0,...\,,N-1,$ (54a)
${\beta}_{q}(\vec{r})$ $=\beta_{q}^{\textrm{Efn}}(\vec{r}),\qquad q=0,...\,,v=N-2,$ (54b)
which solve the Schrödinger equation
$H^{tr}\beta_{q}^{\textrm{Efn}}(\vec{r})=E_{q}\beta_{q}^{\textrm{Efn}}(\vec{r})$ (55)
@@
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{n=0}^{\mathcal{\color[rgb]{0,0,0}N}-1}a^{\textrm{Efn}}_{qn}\phi_{n}(\vec{r}).$ (56)
In this case, $\mathcal{\color[rgb]{0,0,0}N}$ coefficients $g_{n}(E)=\sum_{q=0}^{N-2}b_{q}(E)\,a^{\textrm{Efn}}_{qn}$ in Eq. (52) are expressed through ${v=N-1}$ unknown energy-dependent coefficients $b_{q}(E)$ in the wave function expansion (1) which should be found within the approach.

### v2

@@ II.2 Formal solutions
and $c$ given by Eq. (32) are $\mathbb{\color[rgb]{0,0,0}N}$-dimensional vectors. We note that our SRF selection indicates which SRF $\bar{\beta}_{N-i-w+1}(\vec{r})$ to remove from the $i>1$ channel. Therefore, changing the choices of which functions to remove from the $i>1$ calculations amounts to replacing one row of $A^{2},A^{3},...,A^{w}$ in Eq. (35) by another. See Section IV and Ref. (Supplemental, , Section III) for the two-channel case.
Equations (33) can be resolved step-by-step. First, we solve this matrix equation for the first entrance channel with $i=1$ and obtain $v$ coefficients $b_{q}^{1}(E)={c}^{1}_{q}$, q=0,,v1q=0,...\,,v-1, and $w$ $S$-matrix elements $\mathbb{S}_{j1}(E)=-c^{1}_{j+v-1}$}, j=1,,wj=1,...\,,w at any desired energy $E$. Next, we solve Eq. (33) for $i=2$. The $S$-matrix element $\mathbb{S}_{12}(E)=\mathbb{S}_{21}(E)=-c^{1}_{v+1}$ in the right-hand side of Eq. (34d) was already obtained at the previous step. With the help of Eq. (33) we now obtain $v$ coefficients $b_{q}^{2}(E)={c}^{2}_{q}$}, q=0,,v1q=0,...\,,v-1, and w1w-1 $S$-matrix elements $\mathbb{S}_{j2}(E)=-c^{2}_{j+v-2}$}, j=2,,wj=2,...\,,w. Now we have $S$-matrix elements $\mathbb{S}_{3j}(E)=-c^{j}_{v+3-j}$ with $j=1$, 2 in the right-hand side of Eq. (33) for $i=3$ and solve it. Next we solve Eq. (33) for $i=4$, 5, etc.
@@
Following the ideas of the HORSE approach Zaitsev1998 ; BangHORSE , the potential energy operator $V$ can be approximated by a finite matrix $V^{tr}$ in the oscillator basis,
$V\approx V^{tr}=\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle\langle\phi_{ni}|V|\phi_{n^{\prime}j}\rangle\langle\phi_{n^{\prime}j}|\\ \equiv\sum_{i,j=1}^{w}\sum_{n=0}^{\mathcal{N}_{i}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j}-1}|\phi_{ni}\rangle V_{ni,n^{\prime}j}\langle\phi_{n^{\prime}j}|,$ (40)
since the potential energy matrix elements in the oscillator basis $V_{ni,n^{\prime}j}$ decrease with $n$ and/or $n^{\prime}$ and can be neglected at large $n$ and/or $n^{\prime}$ as compared to the nonzero kinetic energy matrix elements $T_{ni,n^{\prime}j}=\langle\phi_{ni}|T|\phi_{n^{\prime}j}\rangle=T^{l_{i}}_{nn^{\prime}}\delta_{ij}$,
@@
$-\sum_{j^{\prime}=1}^{i-1}\sum_{j=1}^{w}\sum_{n=0}^{\bar{\cal M}_{j}-1}\sum_{n^{\prime}=0}^{\mathcal{N}_{j^{\prime}}-1}\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}C^{+}_{n^{\prime}l_{j^{\prime}}}\!(k)\,V_{nj,n^{\prime}j^{\prime}}c^{j^{\prime}}_{v+i-j^{\prime}}-\sum_{j^{\prime}=1}^{i-1}\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{kS_{0l_{j^{\prime}}}(k)}c^{j^{\prime}}_{v+i-j^{\prime}},\hskip-7.74998pt\quad$ $\begin{cases}i>1,\\ q=0,...\,,N-i,\end{cases}$ (43d)
@@
where $\bar{\mathcal{M}}_{i}=\min\{\bar{M}_{i},\mathcal{N}_{i}\}$, $\big{(}\bar{a}_{qn;i}^{j}\big{)}^{\!*}$ is a complex conjugate of $\bar{a}_{qn;i}^{j}$, and, as previously mentioned, we set $\bar{a}_{qn;i}^{j}=a_{qn}^{j}$ for q=0,,v1q=0,...,{v-1} and arbitrarily chosen coefficients for q=v,,Niq=v,...,{N-i}.
@@
Efros mentioned VDEfros that, contrary to the conventional Hulthén–Kohn variational method Hulthen1948 ; Kohn1948 ; Moses1953 ; Moses , his approach does not make use of “free-free” matrix elements like $\langle\eta^{\pm}_{i}|H-E|\eta^{\pm}_{j}\rangle$ but, instead, includes only “bound-bound” $\langle\bar{\beta}^{i}_{q}|{H-E}|\beta_{q^{\prime}}\rangle$ and “bound-free” $\langle\bar{\beta}^{i}_{q}|{H-E}|\eta^{\pm}_{j}\rangle$ matrix elements which are easier to calculate. We note that Eqs. (43) include matrix elements of the “bound-bound” type only, which are evaluated within a restricted Hilbert space spanned by a limited set of oscillator functions.
@@
This is significant because the evaluation of the matrix elements $\langle\bar{\beta}^{i}_{q}|{H-E}|\eta^{\pm}_{j}\rangle$ as “bound-bound” (i. e., within a restricted Hilbert space) facilitates calculations. In particular, as seen in Eqs. (43b)–(43d), the term $\langle\bar{\beta}_{q}^{i}|T-E|\eta_{j}^{\pm}\rangle$ disappears due to Eq. (39) and coupling to continuum is represented by $\pm\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{k\,S_{0l_{j^{\prime}}}\!(k)}$. Because the remaining terms that include the interaction and coupling to continuum are in the restricted Hilbert space, we obtain a convenient representation of the matrix elements $\langle\bar{\beta}_{q}^{i}|H-E|\eta_{j}^{\pm}\rangle$. This is especially useful for many-body problems.  We should note however, that in many-body applications, Eqs. (43b)–(43d) become more complicated. The problem is that the terms $\pm\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j^{\prime}}\big{)}^{\!*}}{k\,S_{0l_{j^{\prime}}}\!(k)}$ in these equations account for the kinetic energy of relative motion of colliding particles $T^{rel}$ only. In the many-body case, the colliding particles have an internal structure. Therefore, the Hamiltonian includes the kinetic energy of the internal motion of these particles $T^{int}$, which enters Eqs. (43b)–(43d) in the same manner as potential energy terms. The separation of the total kinetic energy operator $T=T^{rel}+T^{int}$ into the relative $T^{rel}$ and internal $T^{int}$ parts may be nontrivial within NCSM but will be much more natural and simpler in approaches like the RGM.
@@
In the case of the standing-wave approach starting from the wave-function ansatz (19), one can use all the above equations of this subsection with replacements $\mathbb{S}_{ij}(E)\rightarrow K_{ij}(E)$, $\tilde{h}^{-}_{l_{i}}(kr)\rightarrow j_{l_{i}}(kr)$, $\tilde{h}^{+}_{l_{i}}(kr)\rightarrow\tilde{n}_{l_{i}}(kr)$}, $C^{-}_{nl_{i}}(k)\rightarrow S_{nl_{i}}(k)$, $C^{+}_{nl_{i}}(k)\rightarrow C_{nl_{i}}(k)$}, having in mind that with these replacements, Eqs. (37) should be substituted by the following relations:
[…]
@@
As a result, Eqs. (43a)–(43b) preserve their structure with the replacement $C^{+}_{nl}(k)\rightarrow C_{nl}(k)$ while in Eqs. (43c)–(43d) the substitution $C^{-}_{nl}(k)\rightarrow S_{nl}(k)$ should be accompanied by removing the terms $\frac{\hbar^{2}}{2m}\frac{\big{(}\bar{a}_{q0;i}^{j}\big{)}^{\!*}}{kS_{0l_{j}}\!(k)}$.
@@ II.3 $S$-matrix poles
The vector $B^{i}$ at $i>1$ includes $S$-matrix elements [the components $c^{j}_{v+i-j}$ of the vector $c$, see Eqs. (34d) and (31)] which have poles; thus searching for poles of the vector $c^{i}$ by locating the zeros of $\det A^{i}$ with $i>1$ is not justified and one can omit some of the $S$-matrix poles. Locating the zeros of $\det A^{1}(E)$ in the complex energy plane can be efficiently performed by the technique of calculating contour integrals described in Ref. AMShirokovHORSE2018 . Other techniques for locating $S$-matrix poles are presented in Refs. Moiseyev1998 ; Fanto2018 .
@@
Note, Eq. (46) guarantees that Eq. (48) has a nonzero solution. However, the vector of unknowns $c^{1}$ is obtained up to an arbitrary multiplier, and the bound state wave function $u_{i}(E;\vec{r})$ should be additionally normalized. Moreover, the bound state wave function in this case is expressed as an infinite converging series of oscillator functions since the function $\eta^{+}_{i}(\vec{r})$ entering $u_{i}(E;\vec{r})$ is defined through infinite oscillator function series; see Eqs. (7) and (15).
@@ III Single-Channel Scattering
We proceed to discuss the efficiency of the adapted Efros approach and its advantages and disadvantages using two-body single-channel scattering. As the first example, we consider $np$ scattering in the $^{1}S_{0}$} channel using the Daejeon16 $NN$ interaction Daejeon16 . The Daejeon16 interaction is defined as a finite matrix and therefore, the HORSE method gives the exact solution that can be used for comparison with our calculated results. In the single-channel case, it is convenient to utilize the standing-wave approach with real-valued wave functions where the only $K$-matrix element in Eq. (19) is expressed through the scattering phase shifts $\delta_{l}(E)$:
@@
$H^{tr}=\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle\langle\phi_{n}|H|\phi_{n^{\prime}}\rangle\langle\phi_{n^{\prime}}|\equiv\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle H_{nn^{\prime}}\langle\phi_{n^{\prime}}|,$ (50)
where $\mathcal{N}$ is the number of harmonic oscillator basis functions used in the truncated wave function expansion. The truncation corresponds conventionally to the maximal value of (excitation) oscillator quanta $N_{\max}$. If we naively suppose that the Hamiltonian structure conventionally employed in ab initio NCSM calculations is the Hamiltonian of the system (i.e., we set $H=H^{tr}$), then we get incorrect results as $H^{tr}$ does not include coupling to continuum. That is discussed in Ref.  (Supplemental, , Section I).
In the case of single-channel two-body scattering, the relative motion oscillator functions $\phi_{n}(\vec{r})$ are given by Eq. (3) and $2n+l\leq N_{\textrm{max}}$, where $n=0,1,...\,,\mathcal{N}-1$. Equation (12) takes the form
@@
[…]
$\hskip-5.59721pta^{as}_{n}(E)$ $=S_{nl}(k)+C_{nl}(k)\tan\delta_{l}(E),\quad n=0,$$\infty$, (53)
and $\mathcal{N}$ energy-dependent coefficients $g_{n}(E)$ should be found within the approach. In our case of $NN$ scattering in the $^{1}S_{0}$ partial wave, the fixed quantum numbers are $l=0$, $S=0$, $J=0$, and $M_{J}=0$; therefore $\mathcal{N}=\frac{1}{2}N_{\max}+1$}.
The asymptotic expansion coefficients $a^{as}_{n}(E)$ describing the free motion at energy $E$ and within the HORSE formalism fit the free Schrödinger equation Zaitsev1998 ; BangHORSE
$\sum_{n^{\prime}=0}^{\infty}(T^{l}_{nn^{\prime}}-E\delta_{nn^{\prime}\!})\,a^{as}_{n^{\prime}}(E)=0,\qquad n=1,$$\,\infty$, (54)
Therefore, according to our proposition regarding the structure of the wave function expansion coefficients, Eqs. (52) and (53), the Hamiltonian (50) should be extended to infinity in the oscillator space by the kinetic energy matrix elements,
$H=\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle V_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=0}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ \equiv\sum_{n,n^{\prime}=0}^{\mathcal{N}-1}|\phi_{n}\rangle H^{tr}_{nn^{\prime}}\langle\phi_{n^{\prime}}|+\sum_{n,n^{\prime}=\mathcal{N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|\\ +|\phi_{\mathcal{N}-1}\rangle T^{l}_{\mathcal{N}-1,\mathcal{N}}\langle\phi_{\mathcal{N}}|+|\phi_{\mathcal{N}}\rangle T^{l}_{\mathcal{N},\mathcal{N}-1}\langle\phi_{\mathcal{N}-1}|,$ (56)
where the truncated Hamiltonian $H^{tr}$ given by Eq. (50) has the structure conventionally used in variational calculations like NCSM and its matrix elements $H^{tr}_{nn^{\prime}}$ include both potential energy matrix elements $V_{nn^{\prime}}$ and kinetic energy matrix elements $T^{l}_{nn^{\prime}}$ with $n,n^{\prime}=0$, … , $\mathcal{N}-1$, the term $\sum_{n,n^{\prime}=\mathcal{N}}^{\infty}|\phi_{n}\rangle T^{l}_{nn^{\prime}}\langle\phi_{n^{\prime}}|$ describes the free motion (continuum) of colliding particles, and the last two terms in Eq. (56) describe the coupling between the continuum and the interaction region in the oscillator space. The role of these interaction region–continuum coupling terms is very important and cannot be neglected. So, for the correct application of the Efros formalism with oscillator basis, one should use in the single-channel case the Hamiltonian (56).
We need to specify the sets of SRFs $\{\beta_{q}\}$ and $\{\bar{\beta}_{q}\}\equiv\{\bar{\beta}^{1}_{q}\}$} entering Eqs. (34). Regarding the SRFs $\beta_{q}(\vec{r})$, one can either use a set of N1N-1 oscillator functions $\phi_{n}(\vec{r})$ with $n\leq\mathcal{N}-1$, eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the Hamiltonian $H^{tr}$, or a set of N1{N-1} linearly independent superpositions of these functions which may be nonorthogonal. As for the functions $\bar{\beta}_{q}(\vec{r})$, we suggest to use N1{N-1} functions $\beta_{q}(\vec{r})$ plus an additional function $\bar{\beta}_{N-1}(\vec{r})$ that guarantees the coupling to the continuum. It is most convenient to set it equal to $\phi_{\mathcal{N}}(\vec{r})$ which is coupled to a single function $\phi_{\mathcal{N}-1}(\vec{r})$ in the interaction region by the analytic matrix element $T^{l}_{\mathcal{N}-1,\mathcal{N}}$. That is, we choose
$\bar{\beta}_{q}(\vec{r})=\ \begin{cases}\beta_{q}(\vec{r}),&0\leq q\leq N-2,\\ \phi_{\mathcal{N}}(\vec{r}),&q=N-1.\end{cases}$ (57)
A natural option is to use $\color[rgb]{0,0,0}N-1\color[rgb]{0,0,0}\leq\mathcal{N}$ eigenfunctions $\beta_{q}^{\textrm{Efn}}(\vec{r})$ of the truncated Hamiltonian (50), which solve the Schrödinger equation
$H^{tr}\beta_{q}^{\textrm{Efn}}(\vec{r})=E_{q}\beta_{q}^{\textrm{Efn}}(\vec{r}),$ (58)
@@
$\beta_{q}^{\textrm{Efn}}(\vec{r})=\sum_{n=0}^{\mathcal{N}-1}a^{\textrm{Efn}}_{qn}\phi_{n}(\vec{r}).$ (59)
In this case, $\mathcal{N}$ coefficients $g_{n}(E)=\sum_{q=0}^{N-2}b_{q}(E)\,a^{\textrm{Efn}}_{qn}$ in Eq. (52) are expressed through v=N1{v=N-1} unknown energy-dependent coefficients $b_{q}(E)$ in the wave function expansion (1) which should be found within the approach.